OpenStudy (anonymous):

A rectangle with a perimeter of 102 inches is half as wide as it is long. Write and solve an equation to determine the length of this rectangle in inches.

OpenStudy (texaschic101):

P = 2L + 2W P = 102 W = 1/2L (half as wide as it is long means its width is half of its length) Now we sub 102 = 2L + 2(1/2L) 102 = 2L + 1L 102 = 3L 102/3 = L 34 = L W = 1/2L W = 1/2(34) W = 17 check.. P = 2L + 2W 102 = 2(34) + 2(17) 102 = 68 + 34 102 = 102 (correct)

OpenStudy (anonymous):

Thank you, I understand now.